# TopoCellGen: Generating Histopathology Cell Topology with a Diffusion Model

### v1

@@ Abstract
Accurately modeling multi-class cell topology is crucial in digital pathology, as it provides critical insights into tissue structure and pathology. The synthetic generation of cell topology enables realistic simulations of complex tissue environments, enhances downstream tasks by augmenting training data, aligns more closely with pathologists’ domain knowledge, and offers new opportunities for controlling and generalizing the tumor microenvironment. In this paper, we propose a novel approach that integrates topological constraints into a diffusion model to improve the generation of realistic, contextually accurate cell topologies. Our method refines the simulation of cell distributions and interactions, increasing the precision and interpretability of results in downstream tasks such as cell detection and classification. To assess the topological fidelity of generated layouts, we introduce a new metric, Topological Fréchet Distance (TopoFD), which overcomes the limitations of traditional metrics like FID in evaluating topological structure. Experimental results demonstrate the effectiveness of our approach in generating multi-class cell layouts that capture intricate topological relationships.
@@ 1 Introduction
Deep-learning methods have made substantial advances in foundational tasks for nuclei analysis, including instance segmentation [29, 21, 40, 32], classification, and detection [1, 52]. These tasks provide the basis for downstream analyses, enabling detailed characterization of tissue architecture and cellular interactions, which are crucial for diagnostic and prognostic applications in pathology [46, 12]. However, accurately annotating multi-class cell arrangements in pathology images is challenging, as distinct cell types display unique spatial patterns, requiring significant domain expertise. Although annotated datasets for multiple cell types exist, they often lack the diversity needed for generalization across various tissues and organs.
To alleviate the burden of manual annotation and enhance the efficiency of analysis, there has been growing interest in utilizing generative models. Early works use Generative Adversarial Networks (GANs) [20] for automatic generation of pathology images [2, 34, 10]. In recent years, diffusion models [61, 31, 50, 11, 73, 74, 45] have emerged as much more reliable alternatives, generating accurate, high-resolution histopathology images [71, 51, 48, 23, 4]. However, all these diffusion models are only trained to directly generate the histopathology images. Despite the impressive visual results, these powerful models deliver limited insight into the underlying biology. It is very hard to connect the learned distributions with human knowledge about tumor microenvionment. This makes it challenging to validate, generalize, or control these models.
@@
Figure 1: Illustrations of intra-class distribution and the inter-class relationship across various cell types. Here we only highlight the tumor/epithelial and lymphocytes.
@@ 4.1 Experimental Results
Qualitative Results. The qualitative results in Fig. 5 illustrate that our proposed method, TopoCellGen, successfully generates cell layouts that closely align with the reference layouts in both cell density and spatial arrangement. In each example, the generated layouts demonstrate a high degree of consistency in cell distribution, preserving the density patterns observed in the reference cell layouts. Furthermore, the spatial relationships between cell types are maintained, with distinct inter-class interactions that reflect the structural characteristics of the reference layouts.
Quantitative Results on Sample Quality.  Tab. 1 shows the results of sample quality and generation accuracy on the BRCA-M2C and Lizard datasets compared with the state-of-the-art cell layout generation baselines: ADM  [11], TMCCG  [2] and Spatial Diffusion  [44]. Across all multi-class datasets, our proposed TopoCellGen demonstrates superior performance by achieving the lowest FID and TopoFD. This means that the generated samples of TopoCellGen have the highest visual fidelity and topological accuracy, closely approximating real multi-class cell distributions. Beyond this, our TopoCellGen also greatly reduces the count errors across multiple cell classes, resulting in more accurate and topologically consistent synthetic cell layouts compared to other methods. These results emphasize the capability of TopoCellGen to maintain realistic cell distribution and topological relationships, meanwhile achieving precise control of cell density.
@@ 4.2 Ablation Studies
Ablation Study on Loss Components. We evaluate the contributions of three loss functions to the model’s performance in generating synthetic cell layouts, measured by FID, Total Counting Error, and TopoFD. As shown in Tab. 3, the results demonstrate that $\mathcal{L}_{\text{count}}$ significantly reduces the counting errors, achieving better accuracy when combined with the other losses. The $\mathcal{L}_{\text{intra}}$ improves the fidelity and topological accuracy of the generated multi-class layouts, as indicated by lower FID and TopoFD. The $\mathcal{L}_{\text{inter}}$ has a smaller effect on the cell counting error but enhances structural consistency, as reflected in improved TopoFD. Using all three losses together yields the best results across all metrics, demonstrating their complementary roles in achieving accurate and realistic multi-class cell layouts.
@@ 5 Conclusion
In conclusion, TopoCellGen presents a robust framework for generating realistic cell topologies in digital pathology by integrating 0- and 1-dimensional topological features within a diffusion model. This method addresses significant limitations of existing approaches, including lack of control over cell count and challenges in accurately capturing both intra-class spatial patterns and inter-class interactions. By integrating topological features into the generative process, our model preserves essential structural characteristics, resulting in realistic and contextually accurate cell layouts. Additionally, the proposed TopoFD metric provides a thorough evaluation of topological fidelity, validating our model’s effectiveness in representing complex spatial and topological relationships. Experimental results demonstrate that our TopoCellGen produces synthetic cell layouts that closely mimic real tissue structures and enhances performance in tasks like cell detection and classification, highlighting its potential to enrich training data and support advancements in digital pathology research.

### v2

@@ Abstract
Accurately modeling multi-class cell topology is crucial in digital pathology, as it provides critical insights into tissue structure and pathology. The synthetic generation of cell topology enables realistic simulations of complex tissue environments, enhances downstream tasks by augmenting training data, aligns more closely with pathologists’ domain knowledge, and offers new opportunities for controlling and generalizing the tumor microenvironment. In this paper, we propose a novel approach that integrates topological constraints into a diffusion model to improve the generation of realistic, contextually accurate cell topologies. Our method refines the simulation of cell distributions and interactions, increasing the precision and interpretability of results in downstream tasks such as cell detection and classification. To assess the topological fidelity of generated layouts, we introduce a new metric, Topological Fréchet Distance (TopoFD), which overcomes the limitations of traditional metrics like FID in evaluating topological structure. Experimental results demonstrate the effectiveness of our approach in generating multi-class cell layouts that capture intricate topological relationships. Code is available at https://github.com/Melon-Xu/TopoCellGen.
@@ 1 Introduction
Deep-learning methods have made substantial advances in foundational tasks for nuclei analysis, including instance segmentation [29, 21, 40, 32, 73], classification, and detection [1, 52]. These tasks provide the basis for downstream analyses, enabling detailed characterization of tissue architecture and cellular interactions, which are crucial for diagnostic and prognostic applications in pathology [46, 12]. However, accurately annotating multi-class cell arrangements in pathology images is challenging, as distinct cell types display unique spatial patterns, requiring significant domain expertise. Although annotated datasets for multiple cell types exist, they often lack the diversity needed for generalization across various tissues and organs.
To alleviate the burden of manual annotation and enhance the efficiency of analysis, there has been growing interest in utilizing generative models. Early works use Generative Adversarial Networks (GANs) [20] for automatic generation of pathology images [2, 34, 10]. In recent years, diffusion models [61, 31, 50, 11, 75, 76, 45, 69] have emerged as much more reliable alternatives, generating accurate, high-resolution histopathology images [72, 51, 48, 23, 4]. However, all these diffusion models are only trained to directly generate the histopathology images. Despite the impressive visual results, these powerful models deliver limited insight into the underlying biology. It is very hard to connect the learned distributions with human knowledge about tumor microenvionment. This makes it challenging to validate, generalize, or control these models.
@@
Figure 1: Illustrations of intra-class distribution and the inter-class relationship across various cell types. Here we only highlight the tumor/epithelial, lymphocytes, and stromal cells.
@@ 4 Experiments
Biological Plausibility. To further validate the quality of our generated cell layouts, we conducted an evaluation involving a board-certified pathologist with 7+ years of experience. From a pathologist’s perspective, the synthetic cell layouts are nearly indistinguishable from real ones and accurately capture the defining benign or malignant features of their real counterparts. More details are in the Supplementary.
@@ 4.1 Experimental Results
Qualitative Results. Qualitative results are shown in Fig. 5. Our proposed TopoCellGen method demonstrates cell layouts that closely match the reference layouts in both density and spatial arrangement. The generated layouts exhibit a high degree of distribution consistency, preserving density patterns and spatial relationships among cell types while reflecting the structural characteristics of the reference layouts.
Quantitative Results on Sample Quality.  Tab. 1 shows the results of sample quality and generation accuracy on the BRCA-M2C and Lizard datasets compared with the state-of-the-art cell layout generation baselines: ADM  [11], TMCCG  [2] and Spatial Diffusion  [44]. Across all multi-class datasets, our proposed TopoCellGen attains the lowest FID and TopoFD, indicating superior visual fidelity and topological accuracy. Moreover, TopoCellGen notably reduces multi-class count errors, resulting in more precise and topologically consistent synthetic cell layouts. These findings highlight TopoCellGen’s capacity to preserve realistic cell distributions, maintain topological relationships, and enable accurate control over cell density.
@@ 4.2 Ablation Studies
Ablation Study on Loss Components. We evaluate the contributions of three loss functions to the model’s performance in generating synthetic cell layouts, measured by FID, Total Counting Error, and TopoFD. As shown in Tab. 3, the results demonstrate that $\mathcal{L}_{\text{count}}$ significantly reduces the counting errors, achieving better accuracy when combined with the other losses. The $\mathcal{L}_{\text{intra}}$ improves the fidelity and topological accuracy of the generated multi-class layouts, as indicated by lower FID and TopoFD. The $\mathcal{L}_{\text{inter}}$ has a smaller effect on the cell counting error but enhances structural consistency, as reflected in improved TopoFD. Combining all three losses yields superior results across all metrics, highlighting their complementary contributions to accurate and realistic multi-class cell layouts.
@@ 5 Conclusion
In summary, TopoCellGen presents a robust framework for generating realistic cell topologies in digital pathology. It accurately preserves both intra- and inter-class spatial patterns, ensures cell count control, and achieves high structural fidelity. Experimental results confirm its close approximation of real tissue layouts, thereby enhancing downstream tasks such as cell detection and classification.
Acknowledgment. We thank Dr. Michael L. Miller for evaluating the biological plausibility of our results. This research was partially supported by the National Science Foundation (NSF) grants CCF-2144901, IIS-2123920, IIS-2212046, the National Institute of Health (NIH) grants R01GM148970, R01GM148970-03S1, R21CA258493, R21CA258493-01A1, R01CA297843, and the Stony Brook Trustees Faculty Award.